# Tuning spin-density separation via finite-range interactions: Dimensionality-driven signatures in dynamic structure factors

### v1

@@ Appendix A CAYLEY-HAMILTON THEOREM
$=$ $\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2g^{\prime}n_{0}\right)\pm\sqrt{\left(\frac{\hbar^{2}k^{2}}{2m}\right)^{2}\left[\left(\frac{\hbar^{2}k^{2}}{2m}+2g^{\prime}n_{0}\right)^{2}-\left(\left(\frac{\hbar^{2}k^{2}}{2m}+2n_{0}g^{\prime}\right)^{2}-\left(2n_{0}g^{\prime}_{12}\right)^{2}\right)\right]}}$ (66)
$=$ $\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2g^{\left(0\right)}n_{0}\pm 2g^{\left(0\right)}_{12}n_{0}+\frac{\hbar^{2}k^{2}}{2m}\left[\frac{4mn_{0}\left(g^{\left(2\right)}\pm g^{\left(2\right)}_{12}\right)}{\hbar^{2}}\right]\right)}$ (67)
$=$ $\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}\lambda_{\pm}+2g^{\left(0\right)}n_{0}\pm 2g^{\left(0\right)}_{12}n_{0}\right)}$ (68)
$=$ $g^{\left(0\right)}n_{0}\sqrt{\frac{\hbar^{2}k^{2}}{2mg^{\left(0\right)}n_{0}}\left(\frac{\hbar^{2}k^{2}}{2mg^{\left(0\right)}n_{0}}\lambda_{\pm}+2\pm 2y\right)},$ (69)
@@ Appendix B REMOVE POWER ULTRAVIOLET DIVERGENCE
Appendix C Detailed derivation of Eq. (LABEL:DSF1D1)
@@
$=$ $\text{Im}\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)2}k^{2}\frac{1}{2v_{\rho,\sigma}^{\left(\text{1D}\right)}k}\left(\frac{1}{\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}-\frac{1}{\omega+v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}\right)$ (74)
$\overset{\gamma\rightarrow 0^{+}}{\approx}$ $\text{Im}\frac{\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}{2}\left(\frac{1}{\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k-i\gamma}-\frac{1}{\omega+v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k-i\gamma}\right)$ (75)
=$\overset{\gamma\rightarrow 0^{+}}{=}$ $\frac{\pi\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}{2}\left[\delta\left(\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k\right)-\delta\left(\omega+v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k\right)\right]$ (76)
=$\overset{\omega>0}{=}$ $\frac{\pi\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}{2}\delta\left(\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k\right).$ (77)

### v2

@@ Appendix A CAYLEY-HAMILTON THEOREM
$=$ $\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2g^{\prime}n_{0}\right)\pm\sqrt{\left(\frac{\hbar^{2}k^{2}}{2m}\right)^{2}\left[\left(\frac{\hbar^{2}k^{2}}{2m}+2g^{\prime}n_{0}\right)^{2}-\left(\left(\frac{\hbar^{2}k^{2}}{2m}+2n_{0}g^{\prime}\right)^{2}-\left(2n_{0}g^{\prime}_{12}\right)^{2}\right)\right]}}$
$=$ $\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2g^{\left(0\right)}n_{0}\pm 2g^{\left(0\right)}_{12}n_{0}+\frac{\hbar^{2}k^{2}}{2m}\left[\frac{4mn_{0}\left(g^{\left(2\right)}\pm g^{\left(2\right)}_{12}\right)}{\hbar^{2}}\right]\right)}$
$=$ $\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}\lambda_{\pm}+2g^{\left(0\right)}n_{0}\pm 2g^{\left(0\right)}_{12}n_{0}\right)}$
$=$ $g^{\left(0\right)}n_{0}\sqrt{\frac{\hbar^{2}k^{2}}{2mg^{\left(0\right)}n_{0}}\left(\frac{\hbar^{2}k^{2}}{2mg^{\left(0\right)}n_{0}}\lambda_{\pm}+2\pm 2y\right)},$
@@ Appendix B REMOVE POWER ULTRAVIOLET DIVERGENCE
Appendix C THE HYDRODYNAMIC ACTION OF THE FINITE-RANGE TERM
Considering density fluctuations, we can re-express the Lagrangian density of the terms related to finite-range interactions in the action Eq. (3) as:
$\mathcal{L}_{\text{finite}}$ $=$ $-\frac{g_{2}}{2}\left(n_{0}+\delta n_{1}\right)\nabla^{2}\left(n_{0}+\delta n_{1}\right)-\frac{g_{2}}{2}\left(n_{0}+\delta n_{2}\right)\nabla^{2}\left(n_{0}+\delta n_{2}\right)$ (69)
$-\frac{g_{212}}{2}\left(\left(n_{0}+\delta n_{1}\right)\nabla^{2}\left(n_{0}+\delta n_{2}\right)+\left(n_{0}+\delta n_{2}\right)\nabla^{2}\left(n_{0}+\delta n_{1}\right)\right)$
$=$ $-\frac{g_{2}}{4}\left(\delta n_{\rho}\nabla^{2}\delta n_{\rho}+\delta n_{\rho}\nabla^{2}\delta n_{\sigma}+\delta n_{\sigma}\nabla^{2}\delta n_{\rho}+\delta n_{\sigma}\nabla^{2}\delta n_{\sigma}\right)$
$-\frac{g_{2}}{4}\left(\delta n_{\rho}\nabla^{2}\delta n_{\rho}-\delta n_{\rho}\nabla^{2}\delta n_{\sigma}-\delta n_{\sigma}\nabla^{2}\delta n_{\rho}+\delta n_{\sigma}\nabla^{2}\delta n_{\sigma}\right)$
$-\frac{g_{212}}{4}\left(\delta n_{\rho}\nabla^{2}\delta n_{\rho}-\delta n_{\rho}\nabla^{2}\delta n_{\sigma}+\delta n_{\sigma}\nabla^{2}\delta n_{\rho}-\delta n_{\sigma}\nabla^{2}\delta n_{\sigma}\right)$
$-\frac{g_{212}}{4}\left(\delta n_{\rho}\nabla^{2}\delta n_{\rho}+\delta n_{\rho}\nabla^{2}\delta n_{\sigma}-\delta n_{\sigma}\nabla^{2}\delta n_{\rho}-\delta n_{\sigma}\nabla^{2}\delta n_{\sigma}\right)$
$=$ $-\frac{\left(g_{2}+g_{212}\right)}{2}\delta n_{\rho}\nabla^{2}\delta n_{\rho}-\frac{\left(g_{2}-g_{212}\right)}{2}\delta n_{\sigma}\nabla^{2}\delta n_{\sigma}$
$=$ $-\frac{g_{2\rho}}{2}\delta n_{\rho}\nabla^{2}\delta n_{\rho}-\frac{g_{2\sigma}}{2}\delta n_{\sigma}\nabla^{2}\delta n_{\sigma}.$
Thus, the action can be obtained as:
$S$ $=$ $\frac{g_{2\rho}}{2}k^{2}\delta n_{\rho}\left(k,\omega\right)\delta n_{\rho}\left(-k,-\omega\right)+\frac{g_{2\sigma}}{2}k^{2}\delta n_{\sigma}\left(k,\omega\right)\delta n_{\sigma}\left(-k,-\omega\right).$ (70)
Appendix D DETAILED DERIVATION OF EQ. (LABEL:DSF1D1)
@@ Appendix C THE HYDRODYNAMIC ACTION OF THE FINITE-RANGE TERM
$=$ $\text{Im}\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)2}k^{2}\frac{1}{2v_{\rho,\sigma}^{\left(\text{1D}\right)}k}\left(\frac{1}{\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}-\frac{1}{\omega+v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}\right)$
$\overset{\gamma\rightarrow 0^{+}}{\approx}$ $\text{Im}\frac{\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}{2}\left(\frac{1}{\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k-\mathrm{i}\gamma}-\frac{1}{\omega+v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k-\mathrm{i}\gamma}\right)$
$\overset{\gamma\rightarrow 0^{+}}{=}$ $\frac{\pi\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}{2}\left[\delta\left(\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k\right)-\delta\left(\omega+v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k\right)\right]$
$\overset{\omega>0}{=}$ $\frac{\pi\chi_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k}{2}\delta\left(\omega-v_{\rho\left(\sigma\right)}^{\left(\text{1D}\right)}k\right).$